OpenStudy (rva.raghav):

H12P3:Our specification is that we want there to be two break frequencies for this filter: 12Hz and 17kHz. We also want midband frequencies, such as signals around 1kHz, to be amplified by about a factor of 2. We have available a 1.2μF and a 470pF capacitor. We also have a variety of resistors between 100Ω and 100kΩ. However, our op amp really likes RS=20kΩ. Your job is to fill in the remaining details of this design.Please help me nurali

OpenStudy (nurali):

1. 11052.426 2. 19919.266 3. 20000 4. 1.2μF 5. 470pF 6. 11052.426 7. 19919.266 8. 1.2μF 9. 470pF 10. -1.80

OpenStudy (nurali):

if correct please click best response.

OpenStudy (anonymous):

could you please mention the formula using for calculations

OpenStudy (anonymous):

HWP3 filter: 10Hz and 19kHz. signals around 1kHz, to be amplified by about a factor of 2. 0.68μF and a 390pF capacitor. resistors between 100Ω and 100kΩ. RS=20kΩ. pi=3.14 part1) First find C1=0.68*10^-6 = 6.8e-7 (answer of part 4) Then R1*C1=1/(2*pi*f) R1*C1=1/(2*pi*10) so R1= 23417 part 2)First find C2=390*10^-12 = 3.9e-10 (answer of part 5) R2*C2=1/(2*pi*19*10^3) R2= 21489 part 3) Rs = Rf = 20kΩ = 20000 Ω part 6 : R3= R1 part 7 : R4= R2 part 8: c3=c1 part 9 : c4= c2 part 9 :Midband gain = ­(R4/R3)